please answer its urgent.
evaluate the following integrals, include procedure.

था - do ४) de | + me . १ ) Cose de 3 + sen o
Answer #1

Page Solve the integral 2 IT 1 do 1 + 1 Sino 2. 2T 2 I do 2+ smo 2 t 2. I alssmao = 2 Sino cose 2 + 2 sono coso 2 2 L It smoDate Page 2 IT secco = sanott B 1+ tan % tau a + 2 Now let tau o t - dt seco 2. 2 1 do = dt (utano + 2. 2 do २de it 2+ L t=0Date Page tan (t +/) + 2. 2 tan 2 D . (4+1) 2 Vs V3 Now al change t into o tauo t (o=2 tautt) e 4 tant 2 √3 ( lano + 2) 3 NowPage DIT 97 coso do 3t smo let t 3t suo Cos o do dt dt t log t + c Change (t = 3+Suo) log (3+ smo) + c Now lent o to 21T भा 0